# Foundation Models for Time Series Analysis: A Tutorial and Survey

### v2

@@ Abstract.
Time series analysis stands as a focal point within the data mining community, serving as a cornerstone for extracting valuable insights crucial to a myriad of real-world applications. Recent advancements in Foundation Models (FMs) have fundamentally reshaped the paradigm of model design for time series analysis, boosting various downstream tasks in practice. These innovative approaches often leverage pre-trained or fine-tuned FMs to harness generalized knowledge tailored specifically for time series analysis. In this survey, we aim to furnish a comprehensive and up-to-date overview of FMs for time series analysis. While prior surveys have predominantly focused on either the application or the pipeline aspects of FMs in time series analysis, they have often lacked an in-depth understanding of the underlying mechanisms that elucidate why and how FMs benefit time series analysis. To address this gap, our survey adopts a model-centric classification, delineating various pivotal elements of time-series FMs, including model architectures, pre-training techniques, adaptation methods, and data modalities. Overall, this survey serves to consolidate the latest advancements in FMs pertinent to time series analysis, accentuating their theoretical underpinnings, recent strides in development, and avenues for future research exploration.
Q. Wen is the corresponding author. Email: qingsongedu@gmail.com
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Time series data are characterized by their sequential order and temporal dependencies, encapsulating valuable information about the dynamics of diverse systems and processes (Zhang et al., 2023; Jiang et al., 2024; Wang et al., 2024a). Various time series data (e.g., stock price, traffic flow, electricity) present unique challenges and opportunities for computational analysis, each requiring tailored approaches to effectively capture their inherent properties. The analysis and understanding of time series data is an important piece of data mining, facilitating crucial insights and decisions in many domains (Wen et al., 2022; Jin et al., 2023a), including finance (Yu et al., 2023; Chen et al., 2023f; Mulvey et al., 2022), healthcare (Liu et al., 2023b; Li et al., 2023), cloud computing (Zhang et al., 2021; Yang et al., 2023), environments (Pathak et al., 2022; Chen et al., 2023b), energy (Zhu et al., 2023; Qin et al., 2023), and urban computing (Wang et al., 2023c; Shao et al., 2022).
@@
In recent years, the advancements of deep learning have revolutionized the field of time series analysis. The motivation behind deep learning techniques lies in their ability to automatically learn comprehensive representations from raw data, thus capturing complex nonlinear relationships and temporal dependencies without the need for manual feature engineering. Such capability leads to significant performance improvements compared with traditional statistical methods across numerous time series applications. Foundation models (FMs), such as large language models (LLMs) in natural language processing (NLP) (Zhao et al., 2023) and advanced models in computer vision (CV) (Awais et al., 2023), have emerged as powerful paradigms capable of achieving state-of-the-art performances in their respective fields. The success of these FMs can be attributed to their ability to leverage vast amounts of data to cultivate general-purpose representations, subsequently fine-tuning them, or even deploying them directly in a zero-shot manner to excel across a diverse spectrum of downstream tasks. This approach not only economizes on the need for task-specific model development but also encapsulates a broad understanding of the world, endowing these models with exceptional versatility and efficiency (Bommasani et al., 2021).
@@ 2. Background
Specifically, we look through TSFMs from a methodology perspective: the components of foundation models encompass the data modality, architecture, pre-training, and adaptation technicals: 1) data modality refers to the type of data used for model training, from single modality such as time series, text, images, and audio to multimodality; 2) architecture refers to which deep neural network is adopted as the backbone of the FM, with Transformers (Vaswani et al., 2017; Wen et al., 2023b) being a popular choice for their ability to handle sequential data effectively; 3) Pre-training involves how to train the model on large, diverse datasets to gain a broad understanding of the data, using techniques like supervised learning or self-supervised learning; 4) Adaptation, such as fine-tuning or few-shot learning, is employed to accommodate the pre-trained FMs to specific tasks. This comprehensive framework of FMs, spanning from data modality to adaptation, facilitates the understanding of using them in time series analysis.
@@
Definition 4 (Event Sequence). An event sequence is a temporally ordered set of events that describe the progression of actions or occurrences within a specific context. It can be formalized as =$\mathcal{E}=\{(e_{1},t_{1}),(e_{2},t_{2}),\ldots,(e_{n},t_{n})\}$, where each $e_{i}$ is an event described by a predicate-argument structure that captures the nature of the interaction or occurrence, and $t_{i}$ denotes the timestamp at which the event $e_{i}$ occurs.
@@ 3. Taxonomy
The proposed taxonomy is illustrated in Figure 3, and the related works can be found in Table 2 in the Appendix. The proposed taxonomy unfolds a structured and comprehensive classification to enhance the understanding of foundation models on time series analysis. It is organized into four hierarchical levels, starting with the data category, followed by the model architecture, pre-training techniques, and finally, the application domain. Unlike previous taxonomies, ours distinguishes itself by delving deeper into the foundation models from the methodology perspective, with a keen focus on their architectural designs, pre-training, and adaptation techniques. This method-centric view is pivotal for researchers, providing valuable insights into the mechanisms of why and how foundation models show great potential for time series analysis.
@@
From the methodology perspective: i) regarding model architecture, the proposed taxonomy highlights three primary categories: Transformer-based, non-Transformer-based, and diffusion-based models. Transformer-based models leverage self-attention mechanisms to capture long-range dependencies within time series data, offering significant advantages in handling sequential data. Non-transformer-based models, with their diverse architectures, cater to a wide range of time series tasks by efficiently processing spatial and temporal patterns. Diffusion-based models, a novel addition, employ stochastic processes to model the data generation process, presenting innovative solutions for time series analysis. ii) In terms of pre-training techniques, the proposed taxonomy divides them into fully-supervised and self-supervised methods, the latter of which includes contrastive, generative, and hybrid approaches. This classification shows how different foundation models are trained with or without labels. iii) Adaptation strategies, such as zero-shot learning, prompt engineering, tokenization, and fine-tuning, further exemplify the versatility of foundation models in customizing to specific time series applications.
@@ 4.1. Standard Time Series
Standard time series possess diverse properties, including varying sampling rates and temporal patterns, which pose significant challenges in developing relevant foundation models. These models aim to identify universal patterns within extensive time series data from varied sources, either to enhance specific tasks or for broad time series analysis.
Most of the existing attempts are in the category of task-oriented standard time series foundation models. They leverage single or multiple data modalities to craft robust models targeting particular time series tasks, typically forecasting or classification. For models involved only in a single (i.e., time series) modality, they may either be developed from scratch or on existing pre-trained models from other domains like large language or vision models (Zhao et al., 2023; Du et al., 2022).
@@
Beyond approaches that focus solely on a single data modality of time series, there have been initiatives towards developing multi-modal, task-oriented foundation models. A notable example is Time-LLM (Jin et al., 2023b), which introduces a reprogramming framework to integrate textual and time series information, repurposing an existing LLM into time series forecasters without additional computational costs. In a similar vein, METS (Li et al., 2023) employs a trainable ECG encoder alongside a frozen language model to process paired ECG and clinical reports. Further, there is emerging research on directly prompting LLMs for specific time series tasks. For instance, PromptCast (Xue and Salim, 2022) converts numerical inputs and outputs into prompts, framing the forecasting task as a sentence-to-sentence conversion to leverage language models directly for forecasting. Other studies, such as one involving LLMs prompted with historical stock price data, company metadata, and past economic/financial news, aim to enhance stock return forecasting (Yu et al., 2023). Another example combines a graph neural network with ChatGPT [^1] https://platform.openai.com/docs/api-reference/introduction to predict stock movements (Chen et al., 2023f), illustrating the diverse applications of these methodologies. Additional noteworthy efforts include (Xie et al., 2023) and (Liu et al., 2023b).
Notably, recent efforts have been directed towards creating general-purpose, single-modality standard time series foundation models. TS2Vec (Yue et al., 2022) represents a pioneering effort by introducing a universal framework for learning time series representations through contrastive learning. SimMTM (Dong et al., 2023) explores cross-domain applications, where pre-trained models via masked time series modeling exhibit superior fine-tuning performance in forecasting and classification tasks. More recent works, such as Timer (Liu et al., 2024d) and UniTS (Gao et al., 2024), further advance the field by facilitating general time series analysis through single, large-scale pre-trained models. Moreover, there is a growing interest in adapting pre-trained models, such as LLMs, for broad time series analysis applications. OFA (Zhou et al., 2023b) and TEST (Sun et al., 2023) exemplify this trend, though both approaches necessitate end-to-end fine-tuning for specific tasks.
@@ 5.1.1. Transformer-based Models
The architecture of FMs has seen a significant convergence towards the Transformer (Vaswani et al., 2017), a model architecture first introduced by Vaswani et al. in 2017. The core innovation of the Transformer lies in its utilization of the attention mechanism, which allows the model to dynamically focus on different parts of the input data. The attention function can be succinctly described as $\texttt{Attention}(Q,K,V)=\texttt{Softmax}(QK^{T}/\sqrt{d_{k}})V$, where $Q$, $K$, and $V$ represent the queries, keys, and values matrices respectively, each with dimensions $T\times d_{k}$, and $d_{k}$ serves as a scaling factor to moderate the dot products’ magnitude. It is evident from the formula that the attention mechanism has the capability to learn global, long-range dependencies in data. This distinguishes it from previous architectures, which were often limited by their local receptive fields or dependency windows. Besides, the Transformer’s design is inherently friendly to parallelization, which allows for significant scalability, enabling the processing of large datasets and the construction of models with billions of parameters. Such scalability and efficiency in capturing intricate data patterns have led to the widespread adoption of the Transformer architecture beyond its initial application in natural language processing (NLP) (Devlin et al., 2019) to fields including computer vision (CV) (Dosovitskiy et al., 2020), speech (Hsu et al., 2021), video (Arnab et al., 2021), time series (Table 2) and beyond.
@@
In terms of standard time series analysis, the Transformer architecture leverages its sequence modeling capabilities to capture temporal dynamics. This includes either repurposing pretrained LLMs for time series to leverage their preexisting sequence modeling strengths (Xue and Salim, 2022), or directly using the Transformer architecture as a base for TSFMs, training from scratch to achieve models best suited for the specifics of time series data (Garza and Mergenthaler-Canseco, 2023). Besides, various techniques have been innovated to augment the functionality of Transformer models in time series analysis comprehensively. A common practice in TSFMs segments time series into patches, which can effectively encapsulate local dynamics within input tokens (Zhou et al., 2023b; Chang et al., 2023a; Das et al., 2023; Jin et al., 2023b; Rasul et al., 2023; Nie et al., 2022; Cao et al., 2023; Woo et al., 2024). Another critical design is the normalization layer, where reversible instance normalization (Kim et al., 2021) techniques, standardizing data through instance-specific mean and variance then reverting it at the output layer, have found extensive application across the above models. Moreover, specialized approaches such as multi-resolution analysis, exemplified by Moirai (Woo et al., 2024) through the employment of varying patch sizes, and decomposition strategies, as implemented by TEMPO (Cao et al., 2023) via the separation of complex interactions into the trend, seasonal, and residual components, have been shown to enhance model efficacy substantially.
@@
In addition to conventional time series data, the Transformer architecture has demonstrated efficacy across a diverse array of temporal datasets, such as trajectory and healthcare records, as summarized in Table 2. This expansion highlights the Transformer’s versatile capacity for temporal data analysis.
@@ 5.2.1. Pre-training
Note that the aforementioned pre-training methods typically build the model from scratch and obtain the universal knowledge from data with the same modality (i.e., time series). Nevertheless, recent advancements in time series research have heightened the usage of LLMs (Zhou et al., 2023b; Jin et al., 2023b; Xue and Salim, 2022; Chang et al., 2023a; Cao et al., 2023; Gruver et al., 2023; Yu et al., 2023; Chen et al., 2023f; Xie et al., 2023; Liu et al., 2023b, 2024b, 2024c; Xue et al., 2022; Wang et al., 2023a; Shi et al., 2023; Gunjal and Durrett, 2023), VLMs (Wimmer and Rekabsaz, 2023), and AMs (Yang et al., 2021) that are pre-trained from other data modalities (text sequence, image-text sequence, acoustic signals).
@@ 5.2.2. Adaptation
The adaptation phase tailors the TSFM to specific tasks or datasets, enhancing its performance on those tasks by leveraging the learned generic temporal knowledge. We partition existing methods into four main branches, including direct usage, fine-tuning, prompt engineering, and time series tokenization.
@@ 5.3. Modality
During the pre-training and adaptation of TSFMs, existing methods involve either single or multiple data modalities, where standard time series data, trajectory data, raster data, and text data can be treated as different forms with unique domain perspectives. In this subsection, we review the data modalities that are used in existing TSFMs across different domains.

### v3

@@ Abstract.
Time series analysis stands as a focal point within the data mining community, serving as a cornerstone for extracting valuable insights crucial to a myriad of real-world applications. Recent advances in Foundation Models (FMs) have fundamentally reshaped the paradigm of model design for time series analysis, boosting various downstream tasks in practice. These innovative approaches often leverage pre-trained or fine-tuned FMs to harness generalized knowledge tailored for time series analysis. This survey aims to furnish a comprehensive and up-to-date overview of FMs for time series analysis. While prior surveys have predominantly focused on either application or pipeline aspects of FMs in time series analysis, they have often lacked an in-depth understanding of the underlying mechanisms that elucidate why and how FMs benefit time series analysis. To address this gap, our survey adopts a methodology-centric classification, delineating various pivotal elements of time-series FMs, including model architectures, pre-training techniques, adaptation methods, and data modalities. Overall, this survey serves to consolidate the latest advancements in FMs pertinent to time series analysis, accentuating their theoretical underpinnings, recent strides in development, and avenues for future exploration.
Q. Wen is the corresponding author.
Time series, foundation model, deep learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671451isbn: 979-8-4007-0490-1/24/08ccs: Information systems Spatial-temporal systems
@@ 1. Introduction
Time series data are characterized by their sequential order and temporal dependencies, encapsulating valuable information about the dynamics of diverse systems and processes (Zhang et al., 2023; Jiang et al., 2024; Wang et al., 2024a). Various time series data (e.g., stock price, traffic flow, electricity) present unique challenges and opportunities for computational analysis, each requiring tailored approaches to effectively capture their inherent properties. The analysis and understanding of time series data is an important piece of data mining, facilitating crucial insights and decisions in many domains (Wen et al., 2022; Jin et al., 2023a), including finance (Yu et al., 2023; Chen et al., 2023f), healthcare (Liu et al., 2023b; Li et al., 2023), cloud computing (Zhang et al., 2021), environments (Pathak et al., 2022; Chen et al., 2023b), energy (Zhu et al., 2023), and urban computing (Wang et al., 2023c; Shao et al., 2022).
In recent years, the advancements of deep learning, especially the transformer-based models (Vaswani et al., 2017), have revolutionized the field of time series analysis (Wen et al., 2023b). Following the pioneering work (Li et al., 2019), there has been a surge in research interest exploring the application of transformers for time series analysis (Zhou et al., 2021, 2022; Xu et al., 2021). The motivation behind deep learning and transformers lies in their ability to automatically learn comprehensive representations from raw data, thus capturing complex nonlinear relationships and temporal dependencies without the need for manual feature engineering. Such capability leads to significant performance improvements compared with traditional statistical methods across numerous time series applications. Foundation models (FMs), such as large language models (LLMs) in natural language processing (NLP) (Zhao et al., 2023) and advanced models in computer vision (CV) (Awais et al., 2023), have emerged as powerful paradigms capable of achieving state-of-the-art performances in their respective fields. The success of these FMs can be attributed to their ability to leverage vast amounts of data to cultivate general-purpose representations, subsequently fine-tuning them, or even deploying them directly in a zero-shot manner to excel across a diverse spectrum of downstream tasks. This approach not only economizes on the need for task-specific model development but also encapsulates a broad understanding of the world, endowing these models with exceptional versatility and efficiency (Jin et al., 2024; Bommasani et al., 2021).
@@ 2. Background
Concretely, we investigate TSFMs from a methodology perspective: the components of foundation models encompass the data modality, architecture, pre-training, and adaptation technicals: 1) data modality refers to the type of data used for model training, from single modality such as time series, text, images, and audio to multimodality; 2) architecture refers to which deep neural network is adopted as the backbone of FM, with Transformers (Vaswani et al., 2017; Wen et al., 2023b) being a popular choice for their ability to handle sequential data effectively; 3) Pre-training involves how to train the model on large, diverse datasets to gain a broad understanding of the data, using supervised or self-supervised learning; 4) Adaptation, such as fine-tuning or few-shot learning, is employed to accommodate the pre-trained FMs to specific tasks. This comprehensive framework of FMs, spanning from data modality to adaptation, facilitates the understanding of using them in time series analysis.
@@
Definition 4 (Event Sequence). An event sequence is a temporally ordered set of events that describe the progression of actions or occurrences within a specific context. It can be formalized as =$\mathcal{E}=\{(e_{1},t_{1}),(e_{2},t_{2}),\ldots,(e_{n},t_{n})\}$, where $e_{i}$ is an event described by a predicate-argument structure that captures the nature of the occurrence, and $t_{i}$ denotes the timestamp when $e_{i}$ occurs.
@@ 3. Taxonomy
The proposed taxonomy is illustrated in Figure 3, and the related works can be found in Table 2. The proposed taxonomy unfolds a structured and comprehensive classification to enhance the understanding of foundation models on time series analysis. It is organized into four hierarchical levels, starting with the data category, followed by the model architecture, pre-training techniques, and finally, the application domain. Unlike previous taxonomies, ours distinguishes itself by delving deeper into the foundation models from the methodology perspective, with a keen focus on their architectural designs, pre-training, and adaptation techniques. This method-centric view is pivotal for researchers, providing valuable insights into the mechanisms of why and how foundation models show great potential for time series analysis.
@@
From the methodology perspective: i) regarding model architecture, the proposed taxonomy highlights three primary categories: Transformer-based, non-Transformer-based, and diffusion-based models. Transformer-based models leverage self-attention mechanisms to capture long-range dependencies within time series, offering significant advantages in handling sequential data. Non-transformer-based models, with their diverse architectures, cater to a wide range of time series tasks by efficiently processing temporal patterns. Diffusion-based models, a novel addition, employ stochastic processes to model the data generation process, presenting innovative solutions for time series analysis. ii) In terms of pre-training techniques, the proposed taxonomy divides them into fully-supervised and self-supervised methods, the latter of which includes contrastive, generative, and hybrid approaches. This classification shows how different FMs are trained with or without labels. iii) Adaptation strategies, such as zero-shot learning, prompt engineering, tokenization, and fine-tuning, further exemplify the versatility of FMs in customizing to specific time series applications.
@@ 4.1. Standard Time Series
Standard time series possess diverse properties, including varying sampling rates and temporal patterns, which pose significant challenges in developing relevant FMs. These models aim to identify universal patterns within extensive time series data from varied sources, either to enhance specific tasks or for broad analysis.
Most of the existing attempts are in the category of task-oriented standard time series foundation models. They leverage single or multiple data modalities to craft robust models targeting particular time series tasks, typically forecasting or classification. For models involved only in a single (i.e., time series) modality, they may either be developed from scratch or on existing pre-trained models from other domains like large language or vision models (Zhao et al., 2023).
@@
Beyond approaches that focus solely on a single data modality of time series, there have been initiatives towards developing multi-modal, task-oriented foundation models. A notable example is Time-LLM (Jin et al., 2023b), which introduces a reprogramming framework to integrate textual and time series information, repurposing an existing LLM into time series forecasters without additional computational costs. In a similar vein, METS (Li et al., 2023) employs a trainable ECG encoder alongside a frozen language model to process paired ECG and clinical reports. Further, there is emerging research on directly prompting LLMs for specific time series tasks. For instance, PromptCast (Xue and Salim, 2022) converts numerical inputs and outputs into prompts, framing the forecasting task as a sentence-to-sentence conversion to leverage language models directly for forecasting. Other studies, such as one involving LLMs prompted with historical stock price data, company metadata, and past economic/financial news, aim to enhance stock return forecasting (Yu et al., 2023). Another example combines GNNs with ChatGPT to predict stock movements (Chen et al., 2023f), illustrating the diverse applications of these methodologies. Additional noteworthy efforts include (Xie et al., 2023) and (Liu et al., 2023b).
Notably, recent efforts have been moved towards creating general-purpose, single-modality standard TSFMs. TS2Vec (Yue et al., 2022) is a pioneering effort by introducing a universal framework for representing time series via contrastive learning. SimMTM (Dong et al., 2023) explores cross-domain applications, where pre-trained models via masked time series modeling exhibit superior fine-tuning performance in forecasting and classification tasks. More recent works, such as Timer (Liu et al., 2024d) and UniTS (Gao et al., 2024), further advance the field by facilitating general time series analysis through single, large-scale pre-trained models. Moreover, there is a growing interest in adapting pre-trained models, such as LLMs, for broad time series analysis. OFA (Zhou et al., 2023b) and TEST (Sun et al., 2023) exemplify this trend, though both approaches necessitate end-to-end fine-tuning for specific tasks.
@@ 5.1.1. Transformer-based Models
The architecture of FMs has seen a significant convergence towards the Transformer (Vaswani et al., 2017), a model architecture first introduced for NLP tasks. The core innovation of the Transformer lies in its utilization of the attention mechanism, which allows the model to dynamically focus on different parts of the input data. The attention function can be succinctly described as $\texttt{Attention}(Q,K,V)=\texttt{Softmax}(QK^{T}/\sqrt{d_{k}})V$, where $Q$, $K$, and $V$ represent the queries, keys, and values matrices respectively, each with dimensions $T\times d_{k}$, and $d_{k}$ serves as a scaling factor to moderate the dot products’ magnitude. It is evident from the formula that the attention mechanism has the capability to learn global, long-range dependencies in data. This distinguishes it from previous architectures, which were often limited by their local receptive fields or dependency windows. Besides, the Transformer’s design is inherently friendly to parallelization, which allows for significant scalability, enabling the processing of large datasets and the construction of models with billions of parameters. Such scalability and efficiency in capturing intricate data patterns have led to the widespread adoption of the Transformer architecture beyond its initial application in natural language processing (NLP) (Devlin et al., 2019) to fields including computer vision (CV), speech, video, time series (Table 2) and beyond.
@@
In terms of standard time series analysis, the Transformer architecture leverages its sequence modeling capabilities to capture temporal dynamics. This includes either repurposing pretrained LLMs for time series to leverage their preexisting sequence modeling strengths (Xue and Salim, 2022), or directly using Transformer as a base for TSFMs, training from scratch to achieve models best suited for the specifics of time series data (Garza and Mergenthaler-Canseco, 2023). Besides, various techniques have been innovated to augment the functionality of Transformer models in time series analysis comprehensively. A common practice in TSFMs segments time series into patches, which can effectively encapsulate local dynamics within input tokens (Zhou et al., 2023b; Chang et al., 2023a; Das et al., 2023; Jin et al., 2023b; Rasul et al., 2023; Nie et al., 2022; Cao et al., 2023; Woo et al., 2024; Cheng et al., 2024). Another critical design is the normalization layer, where reversible instance normalization (Kim et al., 2021) techniques, standardizing data through instance-specific mean and variance then reverting it at the output layer, have found extensive application across the above models. Moreover, specialized approaches such as multi-resolution analysis, exemplified by Moirai (Woo et al., 2024) through the employment of varying patch sizes, and decomposition strategies, as implemented by TEMPO (Cao et al., 2023) via the separation of complex interactions into the trend, seasonal, and residual components, have been shown to enhance model efficacy substantially.
@@
In addition to conventional time series data, the Transformer architecture has demonstrated efficacy across a diverse array of temporal datasets, such as trajectory and healthcare records, as summarized in Figure 3. This expansion highlights the Transformer’s versatile capacity for temporal data analysis.
@@ 5.2.1. Pre-training
Note that the above pre-training methods typically build the model from scratch and obtain the universal knowledge from data with the same modality (i.e., time series). Nevertheless, recent advancements in time series research have heightened the usage of LLMs (Zhou et al., 2023b; Jin et al., 2023b; Xue and Salim, 2022; Chang et al., 2023a; Cao et al., 2023; Gruver et al., 2023; Yu et al., 2023; Chen et al., 2023f; Xie et al., 2023; Liu et al., 2023b, 2024b, 2024c; Xue et al., 2022; Wang et al., 2023a; Shi et al., 2023; Gunjal and Durrett, 2023), VLMs (Wimmer and Rekabsaz, 2023), and AMs (Yang et al., 2021) that are pre-trained from other data modalities (text sequence, image-text sequence, acoustic signals).
@@ 5.2.2. Adaptation
The adaptation phase tailors TSFM to specific tasks or datasets, enhancing its performance on those tasks by leveraging the learned generic temporal knowledge. We partition prior methods into four branches, including direct usage, fine-tuning, prompt engineering, and time series tokenization (see Figure 5).
@@ 5.3. Modality
During the pre-training/adaptation of TSFMs, prior methods involve single or multiple data modalities, where standard time series, trajectory, raster, and text can be treated as different forms with unique domain perspectives. In this part, we review the data modalities that are used in existing TSFMs across different domains.